### v2

@@ I-A Background and Related Works
Recently, the performance for various ISAC systems was investigated to achieve concurrent sensing and communication, which were classified into the radar-centric systems [5]-[8], the communication-centric systems [10]-[11], and the joint-design systems [12]-[13]. The authors in [5] investigated beamforming designs for the ISAC systems with multiple communication users for scenarios with point or extended targets. In particular, the Cramér-Rao bound (CRB) of the azimuth angle or the response matrix of the target was minimized by designing the dual-functional beamforming matrix. In [6], a new angular waveform similarity metric was proposed and combined with the integrated main-lobe-to-sidelobe ratio of the transmit beampattern to evaluate the waveform ambiguity properties. The complementary waveform and linear precoder matrix were jointly designed to minimize the weighted sum of the integrated main-lobe-to-sidelobe ratio while considering a predefined signal-to-interference-plus-noise ratio (SINR), power, and peak-to-average-power ratio constraints. In [7], the authors investigated the performance of the ISAC systems with multiple targets and communication users. Both the scenarios without and with prior target knowledge at the base station (BS) were considered, and the multi-target response matrix and the reflection coefficients/angles were estimated, respectively. The CRB of multi-target estimation was minimized while satisfying the constraints of a minimum multicast communication rate and a maximum transmit power. The authors in [8] proposed a new robust beamforming scheme for ISAC systems with different point targets. The maximum CRB of the estimated direction of arrival was minimized by designing the beamforming matrix while ensuring the quality-of-service (QoS) for all the communication users. The authors in [9] proposed an integrated sensing, communication, and computation over-the-air framework to enable simultaneous communication, target sensing (TS), and over-the-air computation. The shared and separated schemes were proposed, and the accuracy of over-the-air computation was optimized by jointly designing the data transmission beamformer and the aggregation beamformer or the data beamformer and the radar sensing beamformer. In [10], the authors proposed a new waveform design based on constructive interference for the ISAC system with multiple communication users. Their results show that the proposed scheme can reduce the transmit power for the given constraint on estimation accuracy or increase the communication signal-to-noise ratio (SNR) with a given power budget. The authors in [11] investigated the SNR-constrained and CRB-constrained joint beamforming and reflection design problems for the reconfigurable intelligent surface (RIS)-aided ISAC systems. The beamforming matrix and RIS reflection matrix were jointly optimized to maximize the sum of communication users’ communication rate while considering the SNR/CRB, the BS’s transmission power, and the RIS’s unit-modulus constraints. In [12], the angle and the reflection coefficient of the target and the CRB of the angle estimation were utilized as the performance metric of the point TS scenario. The authors adopted three performance metrics, the trace, the maximum eigenvalue, and the determinant of the CRB matrix, respectively, to estimate the complete target response matrix with the extended TS scenario. Due to the possibility of mutual interference between the multiplexed signals, there is a trade-off between TS and communication when multi-beam transmitted signals were reused for both sensing and communications. The EE of the ISAC with an extended target and a multi-antenna communication user was investigated in [13].  The energy consumption at the transmitter with the on-off scheme was minimized by jointly designing the transmit covariance matrix and the time allocation while taking the QoS for the communication user and the CRB constraints for the target into account. In [14], two optimization problems were formulated for the full-duplex ISAC systems. The power consumption was minimized and the sum rate was maximized by jointly optimizing the downlink dual-functional transmit signal, the uplink receive beamformers at the BS, and the transmit power at the uplink users. In [15], the joint active and passive beamforming design problem for the RIS-assisted full-duplex ISAC systems with multiple communication users was considered. The sum rate was maximized by jointly optimizing the transmit beamformer, the linear postprocessing filters, the users’ uplink transmission power, and the reflected phase shift of RIS. Ref. [16] investigated the joint receiver design for multiple-input multiple-output ISAC, proposing two design strategies for simultaneous symbol detection and target estimation for the radar signals received by the receiver and the signals from the uplink communication users. In [17], the authors analyzed the impact of the successive interference cancellation order on the perceived rate and communication rate in the uplink of a non-orthogonal multiple access ISAC system.
@@
In ISAC systems, the coexisting strong radar signals can be utilized as inherent jamming signals to fight the external eavesdroppers and improve the transmission security [4]. The authors in [25] investigated the joint designs of secure transmit beamforming for ISAC systems with multiple eavesdroppers. For the scenario with the eavesdropping CSI, the maximum eavesdropping SINR of multiple eavesdroppers was minimized by jointly designing the transmit beamformers of communication and radar systems while satisfying the QoS of LUs’, the SINR constraint of radar target echo, and the transmit power budget. For the scenario without the eavesdropping CSI, an AN-aided transmit beamforming scheme was proposed. The total transmit power of the BS and radar was minimized by jointly designing the transmit beamformer and the covariance matrices of the AN vector at the BS and radar under the SINR constraints of LUs’ communication and radar target echoes. In [26], an unmanned aerial vehicle (UAV) was utilized as the RIS-aided ISAC base station (BS) and a secure transmission scheme was proposed to maximize the average achievable rate or the EE by jointly designing the transmit power allocation, the scheduling of users and targets, the phase shifts at RIS, as well as the trajectory and velocity of the UAV.
@@
Proactive eavesdropping (PE), also called wireless surveillance, is an effective method of supervising unauthorized wireless monitoring in the physical layer. To be specific, legitimate eavesdroppers wiretap the illegitimate wireless communication between transmitters and receivers to maintain public safety. Compared with traditional physical layer security (PLS), the main differences are summarized as follows [27]. 1) Application scenarios: In PLS scenarios, the communication between transmitters and receivers is assumed to be rightful, the eavesdroppers are illegitimate, and various schemes are designed to prevent information leakage to the malicious or potential eavesdroppers [18]-[26]. In PE scenarios, it is presumed that the eavesdroppers are legitimate and approved by government agencies to monitor that the suspicious communication between transmitters and receivers, such as criminals or terrorists who may jeopardize public safety [28]-[30]. Thus, in PLS scenarios, all the efforts are from the transmitters and/or receivers, while all the works are from the view of eavesdroppers in PE scenarios. 2) Prerequisites: Two conditions are assumed in PE scenarios. One is that the illegitimate transmitter has the channel state information of the unauthorized link, and there is an adaptive transmission between the suspicious source and the destination. The other is that the eavesdropper has global CSI and can decode unauthorized wireless information with any small error based on perfect self-interference cancellation (SIC) since the eavesdropper fully knows the transmission scheme of the illegal link as well as the encoding method [30, 31]. There are no similar assumed conditions in PLS scenarios. 3) Performance metric: In PLS scenarios, the fundamental metric is the achievable secrecy rate, defined as the (non-negative) difference between the achievable rate of the communication link ($R_{D}$) and the achievable rate of the eavesdropping link ($R_{E}$). Based on secrecy rate, there are two main performance metrics: average secrecy rate and secrecy outage probability, which denote the statistical average or time average of secrecy rate and the probability that the secrecy rate falls below a target rate [18]-[26]. In PE scenarios, the relationship between $R_{D}$ and $R_{E}$ is crucial. When $R_{E}\geq R_{D}$, surveillance is successful, and $R_{D}$ is called the eavesdropping rate. Otherwise, surveillance fails, and the eavesdropping rate is equal to zero [28]-[32]. For the scenarios for content analysis and event-based monitoring, there are two main performance metrics: average eavesdropping rate (AER) and successful eavesdropping probability (SEP), also called as eavesdropping non-interrupt probability [30], [33]. 4) Motivation: The motivation of PLS is to make the secrecy rate as large as possible while the aim in PE is to make $R_{D}$ as large as possible under the condition $R_{E}\geq R_{D}$ [30]. In particular, when $R_{E}<R_{D}$, the eavesdropper will utilize some scheme (such as transmitting jamming signals) to ensure successful eavesdropping ($R_{E}\geq R_{D}$). In the scenarios with $R_{E}>R_{D}$, the eavesdropper will utilize some scheme (such as forwarding signals for $D$) to make $R_{D}$ as large as possible under the condition $R_{E}\geq R_{D}$ [28], [29]. In summary, PE is an effective technology in the physical layer to maintain public safety. All the schemes utilized at upper layers, such as encryption, are out of the scope of this area.
@@
For the PE scenarios, many works propose outstanding schemes that can be divided into two types [28]. One is spoofing relaying-based PE schemes that forward the suspicious signal to the illegitimate receivers, and the other is jamming-based PE schemes that utilize the jamming signals to control the interference at the illegitimate receivers. The spoofing relaying-based PE scheme is suitable for scenarios wherein the eavesdropping channel is of higher quality than the illegal channel. For example, the authors in Ref. [28] proposed a power splitting scheme wherein the signal received at the monitor with multiple antennas was utilized for eavesdropping and spoofing relaying, respectively. The AER was maximized by jointly designing the power splitting ratios and the precoding matrix at the monitor. With the help of an amplify-and-forward spoofing relay, a UAV was utilized to eavesdrop on multiple suspicious terrestrial links in [29]. The AER was maximized by jointly optimizing the precoder at the relay, the receive combiner at the monitor, and the UAV’s trajectory.
@@
For the scenarios wherein the quality of the illegal channel is better than that of the eavesdropping channel, jamming signals are utilized to achieve successful PE [30]. In particular, in Ref. [32], a full-duplex monitor transmits jamming signals while wiretapping illegitimate messages. The legitimate transmitter cooperated with the monitor to control the interference on the illegitimate receiver. The SEP of the multiple carrier interference network was maximized by jointly optimizing the transmit power of the eavesdropper and legitimate receiver, and an algorithm based on the Lagrange duality method was proposed to solve the formulated problem. In [33], the legitimate transmitter with multiple antennas acted as the proactive eavesdropper and utilized the intended communication signal for a legitimate receiver to affect the illegitimate receiver. The SEP of the system was maximized by jointly designing the transmitting and receiving beamformers at the full-duplex eavesdropper. In Ref. [34], two suspicious communication links were simultaneously wiretapped by a particular full-duplex monitor with multiple antennas. The achievable eavesdropping rate region for the minimum-MSE receivers was investigated and closed-form expressions of the upper and lower boundary points were derived. Moreover, they proposed a time-sharing-based jamming scheme to enlarge this region by designing the transmit covariance matrix while taking the power budget into account. The authors in [35] proposed a cooperative PE scheme where the primary and assistant monitors cooperatively wiretapped multiple suspicious links. The eavesdropping EE was maximized by designing jamming power and the scheme between monitors while taking the energy transfer and load balancing constraints into account. The authors in Ref. [36] proposed a new cooperative PE scheme wherein multiple AF full-duplex relaying monitors simultaneously received and forwarded the suspicious signals to the center monitor. The AER was maximized by jointly optimizing the receive combining vector at the central monitor, the precoders at the relays and the transmit covariance matrix at the jammer. Table I outlines the typical works on ISAC.
@@
[18] Eavesdropper SNR and legitimate user SINR beampattern matching errors/ beampattern gains Received SNR of targets Waveform covariance matrix Multiple users, a target potential Eve
[19] SNR of user Radar echo SINR SINR of radar echo Transmit waveform, receive beamformer Multiple users, a point-like target
[20] Secrecy rate CRB Worst secrecy rate Beamforming vector Multiple users, a single target/Eve
[21] Sum secrecy rate Beampattern gain Sum secrecy rate Transmit beamformer, artificial jamming vector, STAR-RIS’s vectors Multiple users, a single target
[22] SINR of users and Eve Beampattern matching MSE and cross-correlation pattern Weighted sum of beampattern matching errors and cross-correlation patterns Transmit beamformer Multiple users, multiple targets
[23] Achievable secrecy rate Beampattern matching MSE and cross-correlation pattern Weighted sum of beampattern matching errors and cross-correlation patterns Transmit covariance matrix and the sensing/AN covariance matrix Single user, multiple targets
[24] Secrecy rate CRB Weighted sum of CRB and secrecy rate Transmit covariance matrix Multiple users, multiple targets/Eves
[25] Communication SINR and eavesdropping SINR Radar output SINR Eavesdropping SINR/ transmit power BS transmit beamformer, radar transmit beamformer, radar receive filter Multiple eavesdroppers, multiple users
[26] Average achievable rate Received power Average achievable rate Transmit power allocation, users and target scheduling, IRS’s phase shifts, trajectory and velocity of the UAV Multiple users, multiple targets
Our work Eavesdropping rate CRB SNR, CRB Waveform covariance matrix An illegal user, multiple targets

### v3

@@ I-A Background and Related Works
Recently, the performance for various ISAC systems was investigated to achieve concurrent sensing and communication, which were classified into the radar-centric systems [6]-[10], the communication-centric systems [11]-[13], and the joint-design systems [14]-[15]. The authors in [6] investigated beamforming designs for the ISAC systems with multiple communication users for scenarios with point or extended targets. In particular, the Cramér-Rao bound (CRB) of the azimuth angle or the response matrix of the target was minimized by designing the dual-functional beamforming matrix. In [7], a new angular waveform similarity metric was proposed and combined with the integrated main-lobe-to-sidelobe ratio of the transmit beampattern to evaluate the waveform ambiguity properties. The complementary waveform and linear precoder matrix were jointly designed to minimize the weighted sum of the integrated main-lobe-to-sidelobe ratio while considering a predefined signal-to-interference-plus-noise ratio (SINR), power, and peak-to-average-power ratio constraints. In [8], the authors investigated the performance of the ISAC systems with multiple targets and communication users. Both the scenarios without and with prior target knowledge at the base station (BS) were considered, and the multi-target response matrix and the reflection coefficients/angles were estimated, respectively. The CRB of multi-target estimation was minimized while satisfying the constraints of a minimum multicast communication rate and a maximum transmit power. The authors in [9] proposed a new robust beamforming scheme for ISAC systems with different point targets. The maximum CRB of the estimated direction of arrival was minimized by designing the beamforming matrix while ensuring the quality-of-service (QoS) for all the communication users. The authors in [10] proposed an integrated sensing, communication, and computation over-the-air framework to enable simultaneous communication, target sensing (TS), and over-the-air computation. The shared and separated schemes were proposed, and the accuracy of over-the-air computation was optimized by jointly designing the data transmission beamformer and the aggregation beamformer or the data beamformer and the radar sensing beamformer. In [11], the authors proposed a new waveform design based on constructive interference for the ISAC system with multiple communication users. Their results show that the proposed scheme can reduce the transmit power for the given constraint on estimation accuracy or increase the communication signal-to-noise ratio (SNR) with a given power budget. The authors in [12] investigated the SNR-constrained and CRB-constrained joint beamforming and reflection design problems for the reconfigurable intelligent surface (RIS)-aided ISAC systems. The beamforming matrix and RIS reflection matrix were jointly optimized to maximize the sum of communication users’ communication rate while considering the SNR/CRB, the BS’s transmission power, and the RIS’s unit-modulus constraints. In [13], the authors proposed a symbiotic sensing and communication framework composed of BS and passive sensing nodes. It considers both fully digital arrays and hybrid analog-digital (HAD) arrays, and maximized the achievable sum rate through beamforming design under the constraint of the Cramér-Rao Lower Bound (CRLB) for two-dimensional angle-of-arrival estimation. In [14], the angle and the reflection coefficient of the target and the CRB of the angle estimation were utilized as the performance metric of the point TS scenario. The authors adopted three performance metrics, the trace of CRB (TCRB), the maximum eigenvalue of CRB (MECRB), and the determinant of CRB (DCRB), respectively, to estimate the complete target response matrix with the extended TS scenario. Due to the possibility of mutual interference between the multiplexed signals, there is a trade-off between TS and communication when multi-beam transmitted signals were reused for both sensing and communications. The EE of the ISAC with an extended target and a multi-antenna communication user was investigated in [15].  The energy consumption at the transmitter with the on-off scheme was minimized by jointly designing the transmit covariance matrix and the time allocation while taking the QoS for the communication user and the CRB constraints for the target into account. In [16], two optimization problems were formulated for the full-duplex ISAC systems. The power consumption was minimized and the sum rate was maximized by jointly optimizing the downlink dual-functional transmit signal, the uplink receive beamformers at the BS, and the transmit power at the uplink users. In [17], the joint active and passive beamforming design problem for the RIS-assisted full-duplex ISAC systems with multiple communication users was considered. The sum rate was maximized by jointly optimizing the transmit beamformer, the linear postprocessing filters, the users’ uplink transmission power, and the reflected phase shift of RIS. Ref. [18] investigated the joint receiver design for multiple-input multiple-output ISAC, proposing two design strategies for simultaneous symbol detection and target estimation for the radar signals received by the receiver and the signals from the uplink communication users. In [19], the authors analyzed the impact of the successive interference cancellation order on the perceived rate and communication rate in the uplink of a non-orthogonal multiple access ISAC system.
@@
In ISAC systems, the coexisting strong radar signals can be utilized as inherent jamming signals to fight the external eavesdroppers and improve the transmission security [4]. The authors in [27] investigated the joint designs of secure transmit beamforming for ISAC systems with multiple eavesdroppers. For the scenario with the eavesdropping CSI, the maximum eavesdropping SINR of multiple eavesdroppers was minimized by jointly designing the transmit beamformers of communication and radar systems while satisfying the QoS of LUs’, the SINR constraint of radar target echo, and the transmit power budget. For the scenario without the eavesdropping CSI, an AN-aided transmit beamforming scheme was proposed. The total transmit power of the BS and radar was minimized by jointly designing the transmit beamformer and the covariance matrices of the AN vector at the BS and radar under the SINR constraints of LUs’ communication and radar target echoes. In [28], an unmanned aerial vehicle (UAV) was utilized as the RIS-aided ISAC base station (BS) and a secure transmission scheme was proposed to maximize the average achievable rate or the EE by jointly designing the transmit power allocation, the scheduling of users and targets, the phase shifts at RIS, as well as the trajectory and velocity of the UAV. In [29], the authors proposed a physical layer security scheme for the rate splitting multiple access system under the constraint of sensing CRB. For the two scenarios of the eavesdropper’s channel being perfect and the channel being unknown, under the premise of taking into account user fairness, the maximum energy efficiency is achieved by jointly optimizing the beam design.
@@
Proactive eavesdropping (PE), also called wireless surveillance, is an effective method of supervising unauthorized wireless monitoring in the physical layer. To be specific, legitimate eavesdroppers wiretap the illegitimate wireless communication between transmitters and receivers to maintain public safety. Compared with traditional physical layer security (PLS), the main differences are summarized as follows [30]. 1) Application scenarios: In PLS scenarios, the communication between transmitters and receivers is assumed to be rightful, the eavesdroppers are illegitimate, and various schemes are designed to prevent information leakage to the malicious or potential eavesdroppers [20]-[28]. In PE scenarios, it is presumed that the eavesdroppers are legitimate and approved by government agencies to monitor the suspicious communication between transmitters and receivers, such as criminals or terrorists who may jeopardize public safety [31]-[33]. Thus, in PLS scenarios, all the efforts are from the transmitters and/or receivers, while all the works are from the view of eavesdroppers in PE scenarios. 2) Prerequisites: Two conditions are assumed in PE scenarios. One is that the illegitimate transmitter has the CSI of the unauthorized link, and there is an adaptive transmission between the suspicious source and the destination. The other is that the eavesdropper has global CSI and can decode unauthorized wireless information with any small error based on perfect self-interference cancellation (SIC) since the eavesdropper fully knows the transmission scheme of the illegal link as well as the encoding method [33, 34]. There are no similar assumption in PLS scenarios. 3) Performance metric: In PLS scenarios, the fundamental metric is the achievable secrecy rate, defined as the (non-negative) difference between the achievable rate of the communication link ($R_{D}$) and the achievable rate of the eavesdropping link ($R_{E}$). Based on secrecy rate, there are two main performance metrics: average secrecy rate and secrecy outage probability, which denote the statistical average or time average of secrecy rate and the probability that the secrecy rate falls below a target rate [20]-[28]. In PE scenarios, the relationship between $R_{D}$ and $R_{E}$ is crucial. When $R_{E}\geq R_{D}$, surveillance is successful, and $R_{D}$ is called the eavesdropping rate (ER). Otherwise, surveillance fails, and the ER is equal to zero [31]-[35]. For the scenarios for content analysis and event-based monitoring, there are two main performance metrics: average ER (AER) and successful eavesdropping probability (SEP), also called as eavesdropping non-interrupt probability [33], [36]. 4) Motivation: The motivation of PLS is to make the secrecy rate as large as possible while the aim in PE is to make $R_{D}$ as large as possible under the condition $R_{E}\geq R_{D}$ [33]. In particular, when $R_{E}<R_{D}$, the eavesdropper will utilize some scheme (such as transmitting jamming signals) to ensure successful eavesdropping ($R_{E}\geq R_{D}$). In the scenarios with $R_{E}>R_{D}$, the eavesdropper will utilize some scheme (such as forwarding signals for $D$) to make $R_{D}$ as large as possible under the condition $R_{E}\geq R_{D}$ [31], [32]. In summary, PE is an effective technology in the physical layer to maintain public safety. All the schemes utilized at upper layers, such as encryption, are out of the scope of this area.
@@
For the PE scenarios, many works propose outstanding schemes that can be divided into two types [31]. One is spoofing relaying-based PE schemes that forward the suspicious signal to the illegitimate receivers, and the other is jamming-based PE schemes that utilize the jamming signals to control the interference at the illegitimate receivers. The spoofing relaying-based PE scheme is suitable for scenarios wherein the eavesdropping channel is of higher quality than the illegal channel. For example, the authors in Ref. [31] proposed a power splitting scheme wherein the signal received at the monitor with multiple antennas was utilized for eavesdropping and spoofing relaying, respectively. The AER was maximized by jointly designing the power splitting ratios and the precoding matrix at the monitor. With the help of an amplify-and-forward (AF) spoofing relay, a UAV was utilized to eavesdrop on multiple suspicious terrestrial links in [32]. The AER was maximized by jointly optimizing the precoder at the relay, the receive combiner at the monitor, and the UAV’s trajectory.
@@
For the scenarios wherein the quality of the illegal channel is better than that of the eavesdropping channel, jamming signals are utilized to achieve successful PE [33]. In particular, in Ref. [35], a full-duplex monitor transmits jamming signals while wiretapping illegitimate messages. The legitimate transmitter cooperated with the monitor to control the interference on the illegitimate receiver. The SEP of the multiple carrier interference network was maximized by jointly optimizing the transmit power of the eavesdropper and legitimate receiver, and an algorithm based on the Lagrange duality method was proposed to solve the formulated problem. In [36], the legitimate transmitter with multiple antennas acted as the proactive eavesdropper and utilized the intended communication signal for a legitimate receiver to affect the illegitimate receiver. The SEP of the system was maximized by jointly designing the transmitting and receiving beamformers at the full-duplex eavesdropper. In Ref. [37], two suspicious communication links were simultaneously wiretapped by a particular full-duplex monitor with multiple antennas. The achievable ER region for the minimum-MSE receivers was investigated and closed-form expressions of the upper and lower boundary points were derived. Moreover, they proposed a time-sharing-based jamming scheme to enlarge this region by designing the transmit covariance matrix while taking the power budget into account. The authors in [38] proposed a cooperative PE scheme where the primary and assistant monitors cooperatively wiretapped multiple suspicious links. The eavesdropping EE was maximized by designing jamming power and the scheme between monitors while taking the energy transfer and load balancing constraints into account. The authors in Ref. [39] proposed a new cooperative PE scheme wherein multiple AF full-duplex relaying monitors simultaneously received and forwarded the suspicious signals to the center monitor. The AER was maximized by jointly optimizing the receive combining vector at the central monitor, the precoders at the relays and the transmit covariance matrix at the jammer. Table I outlines the typical works on ISAC.
@@
[20]
Eavesdropper SNR
and legitimate user SINR
beampattern matching errors/
beampattern gains
Received SNR of targets Waveform covariance matrix
Multiple users,
a target potential Eve
[21] SNR of user Radar echo SINR SINR of radar echo
Transmit waveform,
receive beamformer
Multiple users,
a point-like target
[22] Secrecy rate CRB
Worst secrecy rate Beamforming vector
Multiple users,
a single target/Eve
[23] Sum secrecy rate Beampattern gain Sum secrecy rate
Transmit beamformer,
artificial jamming vector,
STAR-RIS’s vectors
Multiple users,
a single target
[24] SINR of users and Eve
Beampattern matching MSE
and cross-correlation pattern
Weighted sum of
beampattern matching errors
and cross-correlation patterns
Transmit beamformer
Multiple users,
multiple targets
[25] Achievable secrecy rate
Beampattern matching MSE
and cross-correlation pattern
Weighted sum of
beampattern matching errors
and cross-correlation patterns
Transmit covariance matrix and
the sensing/AN covariance matrix
Single user,
multiple targets
[26] Secrecy rate CRB
Weighted sum of CRB
and secrecy rate
Transmit covariance matrix
Multiple users,
multiple targets/Eves
[27]
Communication SINR
and eavesdropping SINR
Radar output SINR
Eavesdropping SINR/
transmit power
BS transmit beamformer,
radar transmit beamformer,
radar receive filter
Multiple eavesdroppers,
multiple users
[28] Average achievable rate Received power Average achievable rate
Transmit power allocation,
users and target scheduling,
IRS’s phase shifts,
trajectory and velocity of the UAV
Multiple users,
multiple targets
Our work ER CRB SNR, CRB
Waveform covariance matrix
An illegal user,
multiple targets